# Linear Bandits with Partially Observable Features

### v1

@@ 4.2 Doubly Robust Lasso Estimator
where $p\in(1/2,1)$ is coupling probability set by the algorithm. To couple the policy of the actual action $a_{t}$ and the pseudo-action $\widetilde{a}_{t}$, we resample both of them until they match. This coupling yields a lower bound for the observation probability which reduces the variance of the DR pseudo-rewards in Eq. 10. Let $\mathcal{M}_{t}$ denote the event where $\widetilde{a}_{t}=a_{t}$ within a specified number of resamples. For given $\delta^{\prime}\in(0,1)$, we set the number of resamples as $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$ so that event $\mathcal{M}_{t}$ occurs with probability at least $1-\delta^{\prime}/(t+1)^{2}$. Resampling allows the algorithm to explore further to find an action that balances between regret minimization and reward estimation.
This coupling replaces $\epsilon_{t}$ greedy policy with a multinomial distribution $\phi_{1,t},\ldots,\phi_{K,t}$. When we use DR estimation with $\epsilon_{t}$ greedy policy, the inverse probability $\epsilon_{t}^{-1}:=\sqrt{t}$ appears in the pseudo-reward (10), and thus the variance of the pseudo-reward explodes. Therefore, we couple the $\epsilon_{t}$ greedy policy with the multinomial distribution (9) to bound the inverse probability weight $\phi_{a,t}^{-1}=O(K)$.
@@
With the pseudo-actions (coupled with the actual actions), we construct the unbiased pseudo-rewards for all $a\in[K]$,
@@
and note that $\check{\boldsymbol{\mu}}_{t}^{L}$ is the imputation estimator that fills in the missing rewards of unselected arms in round $t$.
For $a\neq\widetilde{a}_{t}$, i.e., an arm $a$ that is not selected in the round $t$, we impute the missing rewards using $\widetilde{\mathbf{x}}_{k}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}$. For $a=\widetilde{a}_{t}$, however, the term $\mathbb{I}(\widetilde{a}_{t}=a)y_{a,t}/\phi_{a,t}$ calibrates the predicted reward to ensure the unbiasedness of the pseudo-rewards for all arms. Given that $\mathbb{E}_{\widetilde{a}_{t}}[\mathbb{I}(\widetilde{a}_{t}=a)]=\mathbb{P}(\widetilde{a}_{t}=a)=\phi_{a,t}$, taking the expectation over $\widetilde{a}_{t}$ on both sides of Eq. 10 gives $\mathbb{E}_{\widetilde{a}_{t}}[\widetilde{Y}_{a,t}]=\mathbb{E}_{t-1}[y_{a,t}]=\widetilde{\mathbf{x}}_{a}^{\top}{{\boldsymbol{\mu}}_{\star}}$ for all $a\in[K]$. Although the estimate $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$ may have high error, it is multiplied by the mean-zero random variable $(1-\mathbb{I}(\widetilde{a}_{t}=a)/\phi_{a,t})$, making the pseudo-rewards robust to the error in $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$. The pseudo-rewards can only be computed when $\widetilde{a}_{t}=a_{t}$, so they are used in rounds when the chosen action $a_{t}$ and the pseudo-action $\widetilde{a}_{t}$ match, indicated by the event $\mathcal{M}_{t}$. Since $\mathcal{M}_{t}$ occurs with high probability, we can compute the pseudo-rewards for almost all rounds.
Our DR Lasso estimator, which incorporates pseudo-rewards for estimation, is defined as follows:
$\begin{split}\widehat{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}\right)^{2}+\frac{10\sigma}{p}\sigma\sqrt{2t\log\frac{2Kt^{2}}{\delta}}\Big{\|}\big{(}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}\big{)}^{1/2}\boldsymbol{\mu}\Big{\|}_{1},\end{split}$ (11)
and the following theorem provides a theoretical guarantee that this estimator converges across all arms after several exploration rounds.
@@
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}(\widehat{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{20\sigma}{p}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{t}}.$ (12)
Although the DR Lasso estimator leverages $K$-dimensional feature vectors, its error bound depends only logarithmically on $K$. Usually, such fast convergence is made possible under classical regularity conditions, such as the compatibility condition and the restrictive minimum eigenvalue condition (Bühlmann and van de Geer, 2011; van de Geer and Bühlmann, 2009). Also, the existing Lasso-based bandit approaches (Bastani and Bayati, 2020; Kim and Paik, 2019; Oh et al., 2021; Ariu et al., 2022; Chakraborty et al., 2023; Lee et al., 2025) impose these conditions directly on the feature vectors. In contrast, our approach does not require this assumption, since our augmented features are orthogonal vectors in $\mathrm{R}(\mathbf{X})^{\perp}$ and their average Gram matrix satisfies $\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\geq\min\{1,\lambda_{\min}(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top})\}$. Thus, the convergence rate has only $\sqrt{\log K}$ rate in terms of $K$.
The consistency is proved by bounding the two components of the error in the pseudo-rewards defined in (10): (i) the noise of the reward and (ii) the error of the imputation estimator $\check{\mu}_{t}$. Since (i) is sub-Gaussian, it can be bounded using martingale inequalities. For (ii), the imputation error $\widetilde{x}_{a}^{\top}(\check{\mu}^{L}_{t}-\mu_{\star})$ is multiplied by the mean-zero random variable $\left(1-\frac{\mathbb{I}(\widetilde{a}_{t}=a)}{\phi_{a,t}}\right)$ and thus it can be bounded by $\|\check{\mu}^{L}_{t}-\mu_{\star}\|_{1}/\sqrt{t}$.
@@
1:INPUT: features $\{\mathbf{x}_{a}:a\in[K]\}$, coupling probability $p\in(1/2,1)$, confidence parameter $\delta>0$.
@@
2:Initialize $\widehat{\boldsymbol{\mu}}_{0}=\mathbf{0}_{K}$, the exploration phase $\mathcal{E}_{t}=\emptyset$ and the exploration factor $C_{\mathrm{e}}:=8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$.

### v2

@@ 4.2 Doubly Robust Lasso Estimator
Moreover, one can show that the resampling succeeds with high probability for each round. Let $\mathcal{M}_{t}$ denote an event that the pseudo-action $\widetilde{a}_{t}$ matches the chosen action $a_{t}$ within a specified number of resamples. For a given $\delta^{\prime}\in(0,1)$, we set the maximum number of resamples to $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$, then $\mathcal{M}_{t}$ occurs with probability at least $1-\delta^{\prime}/(t+1)^{2}$. This resampling allows the algorithm to further explore the action space in order to balance regret minimization with accurate reward estimation.
@@
With the pseudo-actions coupled with the actual actions, we construct unbiased pseudo-rewards for all $a\in[K]$ as follows:
@@
where $\check{\boldsymbol{\mu}}_{t}^{L}$ is the imputation estimator that fills in the missing rewards of unselected arms in round $t$, as defined in Eq. 8.
For $a\neq\widetilde{a}_{t}$, i.e., an arm that is not selected in the round $t$, we impute the missing rewards using $\widetilde{\mathbf{x}}_{k}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}$. For $a=\widetilde{a}_{t}$, on the other hand, the term $\mathbb{I}(\widetilde{a}_{t}=a)y_{a,t}/\phi_{a,t}$ calibrates the predicted reward to ensure the unbiasedness of the pseudo-rewards for all arms. Given that $\mathbb{E}_{\widetilde{a}_{t}}[\mathbb{I}(\widetilde{a}_{t}=a)]=\mathbb{P}(\widetilde{a}_{t}=a)=\phi_{a,t}$, taking the expectation over $\widetilde{a}_{t}$ on both sides of Eq. 10 gives $\mathbb{E}_{\widetilde{a}_{t}}[\widetilde{y}_{a,t}]=\mathbb{E}_{t-1}[y_{a,t}]=\widetilde{\mathbf{x}}_{a}^{\top}{{\boldsymbol{\mu}}_{\star}}$ for all $a\in[K]$. Although the estimate $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$ may have a large error, it is multiplied by the mean-zero random variable $(1-\mathbb{I}(\widetilde{a}_{t}=a)/\phi_{a,t})$, which makes the resulting pseudo-rewards defined in Eq. 10 robust to errors of $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$. The pseudo-rewards can only be computed—and thus can only be used—when the pseudo-action $\widetilde{a}_{t}$ matches the actual action $a_{t}$, that is, when the event $\mathcal{M}_{t}$ occurs. Since $\mathcal{M}_{t}$ occurs with high probability, we are able to compute pseudo-rewards for almost all rounds.
Incorporating the pseudo-rewards, $\widetilde{y}_{a,t}$, into estimation, we define our DR Lasso estimator as follows:
$\begin{split}\widehat{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}\right)^{2}\\ &\qquad+\frac{4\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{2t\log\frac{2Kt^{2}}{\delta}}\|\boldsymbol{\mu}\|_{1},\end{split}$ (11)
and the following theorem provides a theoretical guarantee that this estimator converges to the across all arms after a sufficient number of exploration rounds.
@@
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{8\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}}.$
Although the DR Lasso estimator leverages $K$-dimensional feature vectors, its error bound depends only logarithmically on $K$. Such fast convergence is typically achieved under classical regularity conditions, including the compatibility condition and the restrictive minimum eigenvalue condition (van de Geer & Bühlmann, 2009; Bühlmann & van de Geer, 2011). Existing Lasso-based bandit approaches (Kim & Paik, 2019; Bastani & Bayati, 2020; Oh et al., 2021; Ariu et al., 2022; Chakraborty et al., 2023; Lee et al., 2025) generally impose these conditions directly on the feature vectors. In contrast, our approach does not require such assumptions, as the augmented features are orthogonal vectors lying in $\mathrm{R}(\mathbf{X})^{\perp}$. Moreover, their average Gram matrix satisfies $\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\geq\min\{1,\lambda_{\min}(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top})\}$. Thus, the convergence rate scales only as $\sqrt{\log K}$ with respect to $K$.
The consistency is proved by bounding the two components of the error in the pseudo-rewards defined in Eq. 10: (i) the noise of the reward and (ii) the error of the imputation estimator $\check{\boldsymbol{\mu}}_{t}$. Since (i) is sub-Gaussian, it can be bounded using martingale concentration inequalities. For (ii), the imputation error $\widetilde{\mathbf{x}}_{a}^{\top}(\check{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star})$ is multiplied by the mean-zero random variable $\left(1-{\mathbb{I}(\widetilde{a}_{t}=a)}/{\phi_{a,t}}\right)$ and thus the magnitude of the error can be bounded by $\|\check{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star}\|_{1}/\sqrt{t}$. The complete proof is deferred to Section E.2.
@@
1:  INPUT: Observed features $\{\mathbf{x}_{a}:a\in[K]\}$, coupling probability $p\in(1/2,1)$, confidence parameter $\delta>0$.
@@
2:  Initialize $\widehat{\boldsymbol{\mu}}_{0}=\mathbf{0}_{K}$, exploration phase $\mathcal{E}_{0}=\emptyset$, and exploration factor $C_{\mathrm{e}}:=(8K)^{3}\widetilde{\sigma}^{-2}_{\min}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}$.